- Gemini Home
- Telescopes and Sites
- Science Visitors at Gemini
- Observing With Gemini
- Visiting Instrument Policy
- DSSI Speckle Camera (North)
- TEXES (North)
- Integration Time Calculators
- Adaptive Optics
- Magnitudes and Fluxes
- Near-IR Resources
- Mid-IR Resources
- Observing Condition Constraints
- Performance Monitoring
- SV/Demo Science
- Future Instrumentation
- Queue and Schedules
- Data and Results
Change page style:
Status and Availability
Update May 6, 2013
Fast guide (tip-tilt and focus) with the OIWFS. Note that during the few hours with good seeing we verified that the fast guide with the OIWFS has an acceptable performance under IQ=85%. We have pending the tuning of the AO corrections.
- Full guiding (fast + AO) with PWFS2.
Imaging throughput in all bands
Spectroscopic observations of extended objects through the 3-pix slit, 6-pix slit, and 8-pix slits with the JH and HK grisms.
Spectroscopy of point sources objects through the 3-pix slit, 6-pix slit, and 8-pix slits with the JH and HK grisms.
Measure the flexure between the OIWFS and the focal plane
Perform the usual back to the sky calibrations (i.e., IAA, WCS, SF paramenters)
We have just completed the most recent commissioning run with F2. During most of the run we had bad and very bad seeing in the 1 to 4 arcsec interval in Ks. We had a brief period of good seeing (~0.6 in Ks for a couple of hours on the night of 4/29). Overall, we made progress on the commissioning plan taking full advantage of weather conditions that normally would have been considered "weather loss." Unfortunately, the MOS wheel started to behave erratically during the 5th night of the 6 night run. We were able to continue the commissioning by manually moving the MOS wheel to different positions. In addition, we found that the gate valve baffle, used to minimize the internal background, was slightly misaligned.
We are currently warming up the MOS dewar and check this mechanism to ascertain the nature of this fault and fix it. In parallel, we are also warming up the camera to realign the gate valve baffle in order to carry out the commissioning of faint objects.
Despite the seeing conditions, we performed several commissioning tasks that could be checked out from the commissioning plan.
Update April 2, 2013
A previous uncontrolled cool down of the Hawaii-2 detector while powered has created an excessive number of charge traps leading to excessive hot pixels. In addition, the readout noise has changed. We expect to correct these issues with a warm up of the detector to room temperature.
The MOS wheel mechanism started to stall. We need to warm up the MOS dewar in order to inspect and fix this mechanism.
FLAMINGOS-2 is undergoing an unscheduled thermal cycle. This was prompted by two recent faults:
The expected down time is between two to three weeks. So we still expect to have FLAMINGOS-2 available for the start of its commissioning run at the end of April.
Update March 21, 2013
We have performed an initial assessment of the imaging performance of F2. At the best focus on-axis, we obtain fwhm ~ 0.35 arcsec for pinholes located within a radius < 0.5 arcmin, and fwhm 0.44 arcsec for the region between 2 to 2.5 arcmin. The following plot shows the image quality as a function of radius. We also show, as reference, the limits for the image quality (IQ) constraints of 20%, 70%, and 85%. We are still evaluating the effects of the IQ on spectroscopy mode. Please take these values into consideration for the 13B proposal cycle.
This plot shows the radial profile of the fwhm for a Ks 2-pix pinhole. We also show the limits for the IQ conditions: 20%, 70%, and 85%.
The instrument is now mounted on the flex-rig to undergo further tests. After this, the instrument will be installed on the telescope so that we can start its commissioning in late April 2013.
Update March 1, 2013
Extensive re-alignment work has been done on the instrument. At this time, the excessive coma and astigmatism present in the January images has been almost eliminated. However, the current PSF is not yet optimal and we continue to work improving it. We plan one more lab cold cycle in late March before starting the commissioning in April.
The R3K grism has been re-installed in the instrument. We have extracted and masked the damaged region of the replica transmission grating. Moreover, we are planning to purchase a new R3K grism in the next months.
Update January 30, 2013
The instrument was cooled down in early January 2013 after a 10 month shut down period. During this time, we worked on: modifying the lens holders for the entire camera and the field lens, installing a new field lens, installing a new window cover, installing a new detector board for the current Hawaii-2 detector, re-aligning the OIWFS, improving the thermal shielding of the camera dewar, and performing a first check of the overall optical alignment of the F2 optics (lenses and detector).
All mechanisms worked as expected including the Hawaii-2. The mechanisms were exercised and we even performed a limited flexure test that showed no new issues.
- One issue that we still have to work on is the overall optical alignment of the camera. The images obtained during this first cool down present excessive coma. We are in the process of improving the alignment so that we can perform new tests during the next cool down in February.
We also discovered some damage on the surface of the grating of the resolution 3000 grism (R3K). The grism cell was removed from the instrument in order to examine it and avoid further damage. We have contacted the vendor and the manufacturer so that they can help us to determine the reason for the damage. Then, we will decide what steps to take in order to be able to have a R3K grism available as soon as possible.
Update September 12, 2012
A new cell holder for the Collimator lens (L1) has been received at Cerro Pachon. The thermal properties of this new cell holder will be tested by installing the original, broken lens within it (the lens will be glued). After this baseline test, a set of tests will be carried out with the new holder and the new L1. The new L1 lens arrived at Cerro Pachon in mid-August.
During the shutdown, stresses in other lenses, located in the camera, have been uncovered. These stresses are consistent with point contact on the radial rods due to tilt arising from decentering during cool-down. Two of these lenses are being replaced. In order to minimize the risk of another lens fracture, the way all the camera lenses are held will be modified. The required modifications are minimal and are aimed solely at reducing stress without incurring additional delay. They will be implemented while the new L1 assembly is tested and the OIWFS is realigned in the MOS dewar.
- We remain on schedule to be on the sky no later than mid March 2013.
Update August 9, 2012
As a result of the KP root cause analysis, Gemini has designed a different setup to hold L1. Several thermal cycles were performed on the MOS dewar in order to understand the temperature behaviour of the L1 lens and its holder. For this, we used the original L1 lens holder and a "glued" L1 lens fitted with temperature sensors. Moreover, we contracted Quartus Engineering to perform a FEA of the original L1 lens holder under warm, cold, and thermal changing conditions (i.e., thermal cycles of the MOS dewar needed for mask changes). The new design has been revised by internal and external experts including Quartus Engineering. This new design incorporates the lessons learned from the FEA and the results of the thermal cycles tests. The main feature of the new design is a dramatic reduction on the mechanical stress experienced by the L1 lens under all operational conditions. The new design will recycle the current hardware but change the way the lens is supported radially and axially. This new holder is now ready to go in fabrication.
- A new L1 Calcium Fluoride was received at Gemini South on July 11 and a spare is expected to ship from another vendor by the end of July. Those lenses have their sides polished to decrease the internal stress left over by the fabrication process.
- A new Hawaii-2 fan out board has been installed and tested. This new board includes some improvements designed to minimize the risk of possible delamination and improve the thermal control of the detector.
Update April 9, 2012
Our Kepner-Tregoe root-cause analysis concluded that the fracture was not caused by a single factor but by several possible factors across time. Based on all our observations, measurement of the as-built assembly of the lens and mount, and the workmanship of some critical surfaces in contact, the tensile stress is large, which leaves little margin to any process deviations that could occur over time. Based on this and in order to improve the long term robustness of F2 we have decided to:
- Build a new L1 cell with a stress reducing design.
- Retrofit this design if possible for the other lenses.
- Polish the new L1 lens edge to relieve internal manufacturing stress.
- Review and improve the cool-down and warm up processes to avoid future occurrences of the problem.
Update March 12, 2012
Collimator Lens Crack
On February 3rd, while doing routine tests, we discovered what appeared to be vignetting in the FOV (see below). Further inspection revealed the cause to be a major crack in the large collimator lens ("L1").
Figure 1 (left) Diagram showing the location of the collimator lens. (right) K band image showing the apparent vignetting covering half the FOV. The windmill pattern visible is produced by the window cover shutter.
Figure 2. This is the cracked lens. By the time the instrument had been warmed up, the crack had propagated across the entire diameter.
In order to expedite the return of F2 to its commissioning phase, we are pursuing the following actions:
- Two replacement collimator lenses have been ordered, from two different vendors.
- A new lens holder is being designed. We are consulting with internal and external experts in order to validate this new design.
- We continue our analysis of the root cause of the lens failure. Based on our measurements of the size of all as-built parts and forces of the springs, our preliminary analysis focused on calculating the tensile stresses created on the lens at the contact points of the two radial defining rods (see the figure below from a report by G. Perez at Gemini Observatory).We find that the analytical calculated tensile stress on the glass is over 5000 lb/in2.
The picture above shows the lens inside the lens holder prior to disassemble (compare with the previous diagram). The bottom picture shows the location of the radial defining rod and the crack in the lens.
Because of the uncertainty in the planned schedule to get back on the sky, we are not able to offer F2 for regular use during 2012B. Nonetheless, we are working toward making the instrument available to the science community as soon as possible. If we maintain this initial schedule, we anticipate expanding the System Verification phase to maximize the earliest opportunities for science.
Our planned schedule indicates that the Expected Time1 for F2 to be ready for its recommissioning is October 1, 2012. With the following boundaries:
Optimistic time: July 28, 2012
Most likely time September 10, 2012
Pessimistic time December 5, 2012
1According to PERT (Program Evaluation and Review Technique) the Expected time = ( Optimistic + 4 x Most likely + Pessimistic ) / 6
Please check this page periodically for frequent updates about FLAMINGOS-2.
Last update 03/12/2012